OpenStudy (anonymous):

if the lengthn of a hypotenuse of a right triangle is 10 find the lengths of the other sides when the are is maximum. The anwer is 5square root of 2 X 5 square root of 2

OpenStudy (anonymous):

we know that area of a triangle = 1/2 *base * height = 1/2 * p * b and h^2 = p^2 +b^2 

OpenStudy (anonymous):

then??

OpenStudy (anonymous):